OpenStudy (anonymous):

A music store is offering a Frequent Buyers Club membership. The membership costs $22 per year, and then a member can buy CDs at a reduced price. If a member buys 17 CDs in one year, the cost is $111.25. Determine the cost of each CD for a single member and then write a linear equation to represent the total cost y of a one year membership, if x CDs are purchased.

OpenStudy (akashdeepdeb):

Now this person who buys the 17 CDs is a member. And thus from the $111.25 he paid. A part is for the membership. So Cost of 17 CDs is = Total Cost - cost of membership = $111.25 - $22 = x [find the value] And now if 17 CDs cost = x One CD would cost = x/17 Understood till here? :)

OpenStudy (akashdeepdeb):

@Mr. batman ?

OpenStudy (anonymous):

not really but kind of

OpenStudy (akashdeepdeb):

What did you NOT understand there?

OpenStudy (akashdeepdeb):

Let me try to help you out here! :D

OpenStudy (anonymous):

so it just 111.25 - $22 ?

OpenStudy (anonymous):

I don't get it, is it y=89x-22

OpenStudy (akashdeepdeb):

The question WANTS you to think about the cost. It says, "If a MEMBER buys 17 CDs in one year, the cost is $111.25." So if a member buys it. He has to pay for the MEMBERSHIP charges ALSO. So cost of 17 CDs + Cost of Membership = $111.25 Thus the cost of 17CDs would be = $111.25 - $22 [the membership charge] The cost of 17CDs = $111.25 - $22 = $89.25 Let us take one example now: Say you pay 100 dollars for 10 candies How much then are you paying for ONE? Thus one candy = 100 dollars/10 candies. = 10 dollars. Thus one candy costs 10 dollars. Now let us return back to our example: We have 17 CDs now for 89.25 dollars How much would you pay for ONE? Thus ONE CD costs 89.25 dollars/the 17 CDs. = 5.25 dollars. Got it till here? :D

OpenStudy (anonymous):

wait how do you get 89 ?

OpenStudy (akashdeepdeb):

Thus the cost of 17CDs would be = $111.25 - $22 [the membership charge] The cost of 17CDs = $111.25 - $22 = $89.25

OpenStudy (anonymous):

so how do you get the equation

OpenStudy (akashdeepdeb):

Now you have to generalize the thing Now you know that one CD costs = $5.25 Then if x CDs are purchased, the CDs'll cost you = $(5.25 * x) Cost of one year membership = $22 Total cost = y = cost of membership + cost of x CDs Thus, the linear equation in 2 variables x and y is => y = 22 + 5.25x Understood? :D